# Cost-Free Personalization via Information-Geometric Projection in Bayesian Federated Learning

### v1

@@ 6 Conclusions
In this work, we proposed a personalized BFL method that balances local adaptation with global generalization, while explicitly accounting for uncertainty quantification and model calibration. Our approach combines variational learning with information-geometric tools, namely projection and barycenters, to produce client-specific models that remain robust under data heterogeneity, with minimal additional cost. Through experiments on three benchmark datasets, FashionMNIST, SVHN, and CIFAR-10, we demonstrate that our method consistently achieves a favorable trade-off between performance on local and global data distributions. In global evaluations, our personalized models outperform existing baselines in terms of accuracy, calibration, and uncertainty quantification, while also maintaining highly competitive performance in local data. This contrasts with other methods that tend to overfit or underfit across clients. Furthermore, we show that the global model produced by our approach generalizes relatively better than those of the compared state-of-the-art methods. These empirical results highlight the effectiveness of principled, information-geometric personalization combined with variational learning in federated settings. Future work includes exploring adaptive mechanisms to control the degree of personalization, extending our personalization method to non-parametric BFL, and evaluating its performance under additional heterogeneity scenarios, such as feature and quantity shifts.
@@ Appendix A Proof of Theorem 1
Under Assumption 2, (8) is a convex optimization problem. Therefore, we can apply Lagrangian duality to the projection problem. For a fixed Lagrangian multiplier $\lambda$, the minimization problem becomes
$\min_{p\in\mathcal{M}}\mathcal{L}_{\lambda}(p),\quad\text{where}\quad\mathcal{L}_{\lambda}(p)=D(p||p_{g})+\lambda D(p||p_{k}).$
We can rewrite $\mathcal{L}_{\lambda}(p)$ as
$\qquad\mathcal{L}_{\lambda}(p)=\left(\lambda+1\right)\left(\frac{1}{\lambda+1}D(p||p_{g})+\frac{\lambda}{\lambda+1}D(p||p_{k})\right).$
Let $w_{g}$, $w_{k}$ be defined as in (11). Then, the minimization problem becomes
$\min_{p\in\mathcal{M}}\,w_{g}D(p||p_{g})+w_{k}D(p||p_{k}).$ (12)
By Definition 2, the probability distribution $p$ minimizing the weighted sum in (12) coincides with the $D$-Barycenter. This concludes the proof. ∎
Appendix B IVON Hyperparameters
It is worth noting that training with IVON can be sensitive to hyperparameters, particularly the initialization of the Hessian. To ensure reproducibility, we provide in Table 2 the hyperparameters used for each dataset. Here, $N_{k}$ denotes the size of the training dataset of the $k^{th}$ client.
Table 2: IVON Hyperparameters.
params FashionMNIST SVHN CIFAR-10
initial learning rate 0.1 0.1 0.1
final learning rate 0.01 0.01 0.01
weight decay 2e-4 2e-4 2e-4
batch size 64 64 64
ESS $N_{k}$ $N_{k}$ $N_{k}$
initial hessian ($h_{0}$) 5 2 1
MC sample while training 1 1 1
MC sample while testing 10 10 10
Appendix C Comparison between the different aggregation methods considered
@@
We conduct experiments with various combinations of global update and personalization methods, including EAA, RKLB, and WB for global aggregation, and RKLB and WB for personalization. Across all configurations, we observed consistent trends and similar results. To confirm our observations, we apply the Wilcoxon signed-rank test [15] to compare the performance of different aggregation methods across multiple runs (seeds) and datasets. We compute pairwise Wilcoxon tests between all methods to assess whether their performance differences are statistically significant. This allows us to build a matrix of $p$-values showing, for each pair of methods, whether one consistently outperforms the other. The results of the Wilcoxon signed-rank tests presented in Figure 4 show that, in most cases, the $p$-values are high, indicating no statistically significant difference between the aggregation methods. This suggests that, despite small fluctuations in performance across datasets, the aggregation methods perform similarly overall, and no single aggregation approach consistently outperforms the others in a statistically meaningful way for the different metrics considered. Additionally, to illustrate the variability and stability of the aggregation methods across different datasets, we provide violin plots in Figure 5. These plots show the distribution of global model accuracies across multiple random seeds for each aggregation method on FashionMNIST, SVHN, and CIFAR-10. The shape and spread of the distributions offer insights into both the typical performance (median) and variability (spread) of each method. Based on the violin plots, which show that the WB-based aggregation achieves competitive and consistently stable performance across datasets, we restrict the presentation of the experiments reported in the experimental part to this method for both the global update and personalization steps.
@@
Figure 4: Wilcoxon rank-sum test p-values comparing aggregation methods across all datasets for three evaluation metrics: (a) Accuracy, (b) ECE, and (c) NLL. Lower p-values indicate statistically significant differences between methods. Only the lower triangle of each matrix is shown to avoid redundancy.
@@
Figure 5: Accuracy distributions across seeds for different aggregation methods on: (a) FashionMNIST, (b) SVHN, and (c) CIFAR-10 datasets. Lower spread and higher median indicate better and more stable performance.
Appendix D Beyond Federated Learning: Broader Applicability of the Proposed Method
While the use of Information Geometry in BFL is natural, our method also applies to any optimization over probability spaces rather than point estimates. The method can lend itself to domain adaptation: When a model is trained on source data $A$, and new target data $B$ becomes available, retraining only on $B$ may lead to catastrophic forgetting of the source domain. Our projection framework offers a principled trade-off, retaining prior knowledge while integrating new information. This is further enhanced by the flexibility of our method. In particular, the use of IVON combines the efficiency of deterministic learning with the ability to perform inference and adaptation in the posterior space. However, it is important to note that our approach fundamentally differs from standard domain adaptation methods based on optimal transport [11], which operate on data distributions rather than model posteriors.
Application in Incremental Learning
We design a toy example to simulate an incremental learning scenario involving two sequential tasks:
  • Task A, consisting of the first five classes of MNIST.
  • Task B, corresponding to the remaining five classes.

### v2

@@ 7 Conclusions
In this work, we proposed a personalized Bayesian Federated Learning method that balances local adaptation and global generalization while explicitly accounting for uncertainty quantification and model calibration. Our approach combines variational learning with information-geometric tools, namely projection and barycenters, to produce client-specific models that remain robust under data heterogeneity with minimal additional cost. Through experiments on three benchmark datasets, FashionMNIST, SVHN, and CIFAR-10, we demonstrate that our method consistently achieves a favorable trade-off between performance on local and global data distributions. In global evaluations, our personalized models outperform existing baselines in terms of accuracy, calibration, and uncertainty quantification, while maintaining highly competitive performance on local data. This contrasts with other methods, which often tend to overfit or underfit across clients. Furthermore, we show that the global model produced by our approach generalizes better than those of state-of-the-art alternatives. These empirical results highlight the effectiveness of principled, information-geometric personalization combined with variational learning in federated settings. Future work will explore adaptive mechanisms to control the degree of personalization, extend our personalization framework to nonparametric BFL, and evaluate its performance under additional heterogeneity scenarios, such as feature and quantity shifts, as well as on real-world cross-device FL benchmarks with naturally occurring client partitions.
@@ Appendix A Proof of Theorem 1
In our proof, we directly address the problem through the KKT conditions, leveraging the well-known variational structure of projection problems (see Eq. 5.14, Ekeland and Temam (1999)). This approach allows us to highlight the key concepts and ideas underlying the proof without unnecessary technical overhead. All cited Theorems refer to Ekeland and Temam (1999).
Weak or Strong Duality: We first apply Proposition 5.1 (generalization of Slater’s condition) to show that the projection problem is stable. Subsequently, Proposition 2.2 then implies that it is also normal. Finally, Lemma 5.2 verifies the hypotheses of Proposition 2.1, which yields zero duality gap. Consequently, the max-min and min-max formulations attain the same finite optimal value, establishing strong duality.
Relations between the min-max and max-min solutions: Theorem 5.1 ensures that, for fixed Lagrangian multipliers, if one can identify a primal variable that satisfies the saddle-point condition, then this variable solves the primal problem for an appropriate constraint value. As shown in the proof, this is equivalent to minimizing the Lagrangian with respect to the primal variable while keeping the multipliers fixed. This guarantees that the solution obtained via the barycenter coincides with the solution of the original projection problem.
We restate the problem for clarity. The projection problem is defined as:
$\min_{p\in\mathcal{M}}D(p|p_{g})\quad\text{s.t.}\quad D(p|p_{k})\leq r_{k}.$ (11)
Assumption 2 ensures that $D(\cdot|p_{g})$ and $D(\cdot|p_{k})$ are convex in their first argument. Consequently, the problem is a convex optimization problem with a convex inequality constraint.
Because $p_{k}\in\mathcal{S}_{k}$ and the interior of $\mathcal{S}_{k}$ is nonempty (e.g., any $q$ sufficiently close to $p_{k}$ lies in the interior), the constraint admits a Slater point. Therefore, the KKT conditions are both necessary and sufficient for optimality.
The Lagrangian for problem 11 is:
$\mathcal{L}(p,\lambda)=D(p|p_{g})+\lambda(D(p|p_{k})-r_{k}),\qquad\lambda\geq 0.$ (12)
Let $p^{\star}=p_{g,k}$ be the minimizer of 11. The KKT conditions are:
  1. 1.
    Primal feasability: $D(p^{\star}|p_{k})\leq r_{k}.$
  2. 2.
    Dual feasibility: $\lambda^{\star}\geq 0.$
  3. 3.
    Complementary slackness:
    $\lambda^{\star}\big(D(p^{\star}|p_{k})-r_{k}\big)=0.$ (13)
    Since varying the radius $r_{k}$ parametrizes the geodesic between $p_{g}$ and $p_{k}$, and $p_{g}\notin\mathcal{S}_{k}$ in any nontrivial personalization setting, the solution necessarily lies on the boundary, i.e., $D(p^{\star}|p_{k})=r_{k}$, which implies $\lambda^{\star}>0$.
  4. 4.
    Stationarity:
    $\nabla_{p}D(p^{\star}|p_{g})+\lambda^{\star}\nabla_{p}D(p^{\star}|p_{k})=0.$ (14)
Divide equation 14 by $(1+\lambda^{\star})$:
$\frac{1}{1+\lambda^{\star}}\nabla_{p}D(p^{\star}|p_{g})+\frac{\lambda^{\star}}{1+\lambda^{\star}}\nabla_{p}D(p^{\star}|p_{k})=0.$ (15)
Define normalized nonnegative weights
$w_{g}=\frac{1}{1+\lambda^{\star}},\qquad w_{k}=\frac{\lambda^{\star}}{1+\lambda^{\star}}.$ (16)
Then equation 14 becomes:
$w_{g}\nabla_{p}D(p^{\star}|p_{g})+w_{k}\nabla_{p}D(p^{\star}|p_{k})=0.$ (17)
The barycenter of $\{p_{g},p_{k}\}$ with weights $w_{g},w_{k}$ is defined as the minimizer of:
$\min_{p\in\mathcal{M}}\left(w_{g}D(p|p_{g})+w_{k}D(p|p_{k})\right).$ (5)
Since the objective is convex, its minimizer must satisfy the first-order optimality condition:
$w_{g}\nabla_{p}D(p^{\star}|p_{g})+w_{k}\nabla_{p}D(p^{\star}|p_{k})=0.$ (18)
This is exactly equation 17, obtained from the KKT stationarity condition. Thus, the minimizer of the constrained projection problem 11 coincides with the barycenter:
$p^{\star}=p^{*}_{D}(\{p_{g},p_{k}\},\{w_{g},w_{k}\}).$ (19)
This concludes the proof. ∎
Appendix B Comparison between the Different Aggregation Methods Considered
@@
We conduct experiments with various combinations of global update and personalization methods, including EAA, RKLB, and WB for global aggregation, and RKLB and WB for personalization. Across all configurations, we observe consistent trends and comparable results. To confirm these observations, we apply the Wilcoxon signed-rank test Demšar (2006) to compare the performance of different aggregation methods across multiple runs (random seeds) and datasets. We compute pairwise Wilcoxon tests between all methods to assess whether their performance differences are statistically significant. This yields a matrix of $p$-values indicating, for each pair of methods, whether one consistently outperforms the other. The results of the Wilcoxon signed-rank tests, presented in Figure 5, show that in most cases the $p$-values are high, indicating no statistically significant difference between the aggregation methods. This suggests that, despite small fluctuations in performance across datasets, the aggregation methods perform similarly overall, and no single aggregation approach consistently outperforms the others in a statistically meaningful way across the considered metrics. To further illustrate the variability and stability of the aggregation methods across different datasets, we provide violin plots in Figure 6. These plots show the distribution of global model accuracies across multiple random seeds for each aggregation method on FashionMNIST, SVHN, and CIFAR-10. The shape and spread of the distributions offer insight into both the typical performance (median) and variability (spread) of each method. Based on these results, which show that the WB-based aggregation achieves competitive and consistently stable performance across datasets, we restrict the experiments reported in the main text to this method for both the global update and personalization steps.
@@
Figure 5: Wilcoxon signed-rank test $p$-values comparing aggregation methods across all datasets for three evaluation metrics: (a) accuracy, (b) ECE, and (c) NLL. Lower $p$-values indicate statistically significant differences between methods. Only the lower triangle of each matrix is shown to avoid redundancy.
@@
Figure 6: Accuracy distributions across random seeds for different aggregation methods on: (a) FashionMNIST, (b) SVHN, and (c) CIFAR-10 datasets. A lower spread and higher median indicate better and more stable performance.
Appendix C Effect of $\lambda$ on the Trade-off between Performance on Local and Global Data
To complement Section 5.3 in the main text, we report in Figures 7 and 8 the results illustrating the effect of $\lambda$ on the trade-off between performance on local and global data for the FashionMNIST and SVHN datasets.
Figure 7: Effect of $\lambda$ on the trade-off between performance on local and global data distributions. Results are reported for the FashionMNIST dataset. Notably, $\lambda=0$ corresponds to the global model, whereas $\lambda\to\infty$ corresponds to the local model.
Figure 8: Effect of $\lambda$ on the trade-off between performance on local and global data distributions. Results are reported for the SVHN dataset. Notably, $\lambda=0$ corresponds to the global model, whereas $\lambda\to\infty$ corresponds to the local model.
Appendix D Client-Level Fairness Analysis
Client heterogeneity introduces significant challenges in distributed learning architectures, one of which is ensuring algorithmic fairness. Algorithmic fairness refers to the impartial treatment of individuals or groups in automated decision-making processes, without bias, discrimination, or favoritism based on innate or acquired characteristics Saxena et al. (2019); Mehrabi et al. (2021). In the context of FL, where models are collaboratively trained on data from multiple sources representing diverse populations, overlooking fairness considerations can amplify or perpetuate existing societal biases in the global model, leading to unfair outcomes for certain subgroups. To mitigate this issue, a growing body of research has proposed fairness-aware FL algorithms, including approaches based on personalized FL Tan et al. (2023); Jeong and Kountouris (2023), which tailor models to individual client data distributions. Shi et al. (2023) presents a comprehensive taxonomy of fairness-aware FL strategies, covering key aspects such as client selection, optimization, contribution assessment, and incentive allocation.
D.1 Related Work
Fairness in Federated Learning.
Fairness is a multifaceted concept that spans several disciplines, including the social sciences, law, machine learning, and statistics, each offering distinct perspectives and implicitly different definitions. In the context of machine learning, Mehrabi et al. (2021) provides a comprehensive overview of key fairness notions, categorizing them into three primary types: individual fairness, which seeks to ensure similar outcomes for similar individuals; group fairness, which focuses on equal treatment across predefined demographic groups; and subgroup fairness, which combines elements of both individual- and group-based approaches to better capture fairness across a broader range of population segments.
In FL, fairness extends beyond the behavior of a single predictive model to encompass the equitable treatment of clients participating in the distributed training process. Because clients contribute heterogeneous data and resources, an FL system is considered fair if it avoids systematically privileging or disadvantaging certain participants. To formalize fairness in this context, we draw on two long-standing philosophical traditions, utilitarianism and egalitarianism, which have recently been adapted to the federated setting Zhang et al. (2022a); Hu et al. (2022b); Wang et al. (2021); Gao et al. (2024), yet remain underexplored within the BFL paradigm.
  • Utilitarianism, rooted in the works of Bentham and Mill Bentham (1890); Bentham and Mill (2004) and formalized by Maskin Maskin (1978), evaluates the fairness of a system through the aggregate welfare it produces. In FL, this corresponds to optimizing global utility, for example, the average or mean accuracy across clients.
  • Egalitarianism, inspired by Rawls’ difference principle Rawls (1974; 1999), instead focuses on protecting the worst-off participants. Translated to FL, this notion aligns with max–min fairness, which aims to maximize the minimum (worst-case) performance across clients, typically evaluated through the worst-10% accuracy metric.
Wasserstein Barycenters for Fairness.
In the domain of fair learning, recent work has highlighted the effectiveness of WBs as a unifying framework for promoting fairness in ML while preserving strong predictive performance. Chzhen et al. (2020) investigated fair regression by establishing a direct connection between the demographic parity constraint and the WB problem. They showed that the optimal fair predictor, minimizing the squared error while ensuring independence from sensitive attributes, can be derived as the WB of the conditional distributions corresponding to different sensitive groups. This approach leads to a simple yet powerful post-processing technique that transforms any regression model into a fair one without requiring retraining. Importantly, their method provides robust, distribution-free fairness guarantees, improving fairness metrics with only minimal reductions in predictive accuracy.
Building on this foundation, Gaucher et al. (2023) extended the use of WBs to classification tasks. They demonstrated that demographic parity in classification can be achieved by solving a fair regression problem, followed by appropriate thresholding. Their approach underscores the role of the WB in aligning group-wise distributions, thereby reducing disparities across sensitive attributes. In settings with binary sensitive attributes, the barycenter plays a central role in determining optimal classification thresholds, enabling a favorable trade-off between fairness and performance. Together, these works provide both theoretical justification and practical algorithms for transport-based fairness, offering a cohesive framework applicable to both regression and classification.
Our work connects with these findings by incorporating WBs during the aggregation phase. We extend this line of research by analyzing WBs through the lens of client-level fairness definitions, offering a complementary perspective to the discussions presented in this section.
D.2 Effect of Variational Inference Layers on Client-Level Fairness
In this subsection, we investigate how increasing the number of variational inference layers (i.e., Bayesian layers implemented via variational inference) in the model architecture affects fairness outcomes. To this end, we focus on the BA-BFL setting implemented within the Hybrid Bayesian Deep Learning framework Jamoussi et al. (2024), alongside the standard FedAvg baseline.
It is important to note that both the WB and RKLB aggregation strategies reduce to simple arithmetic averaging in the deterministic limit (i.e., when posterior variances vanish). Consequently, BA-BFL with zero Bayesian layers is equivalent to FedAvg, enabling a unified comparison in Figure 9. To ensure a fair comparison, we restrict our analysis to BA-BFL and FedAvg, deliberately excluding additional mechanisms specific to other Bayesian methods, including our IVON-based approach, since the differing optimizers prevent a direct comparison with FedAvg in this context. This controlled setup allows for a focused examination of how the degree of Bayesian modeling, quantified by the number of variational inference layers, influences the fairness of the global model across different clients. However, our analysis does not reveal a consistent trend across experiments, suggesting that the relationship between the number of Bayesian layers and client-level fairness may depend on other factors, such as task complexity.
Figure 9: Impact of Variational Inference Layers on Fairness.
Appendix E Beyond BFL: Broader Applicability of the Proposed Method
E.1 Merging Variational Foundation Models
Foundation Models (FMs) have emerged as powerful general-purpose learners, capable of adapting to a wide range of downstream tasks after large-scale pretraining. However, as data distributions shift and new domains emerge, keeping these models up to date without retraining from scratch remains a major challenge Bommasani and et al. (2021). Continual pretraining Ke et al. (2023); Shi et al. (2024), fine-tuning Hu et al. (2022a), and model merging McMahan et al. (2017); Matena and Raffel (2022); Daheim et al. (2023); Goddard et al. (2024) offer promising paths forward, enabling FMs to integrate new knowledge while retaining broad generalization. However, updating such large systems at scale faces key challenges: computational constraints, catastrophic forgetting McCloskey and Cohen (1989), and potential misalignment in uncertainty quantification. Addressing these challenges calls for principled and efficient update rules that incorporate domain-specific adaptations into a global FM while preserving statistical rigor and scalability.
Within the spectrum of adaptation strategies, variational approaches provide a rigorous framework for representing model uncertainty, making them particularly suitable in contexts where both reliability and interpretability are critical. We focus on variational FMs whose parameters encode posterior distributions, enabling updates to be expressed as operations on the statistical manifold of distributions. We assume these models are pretrained or fine-tuned using the IVON optimizer Shen et al. (2024), chosen for its ability to match Adam’s computational cost while delivering strong Bayesian performance in large-scale settings. IVON has been shown to pretrain GPT-2 on OpenWebText and ResNet on ImageNet from scratch, as well as to fine-tune large masked language models (e.g., DeBERTaV3). Within this setting, we formulate FM merging as an information-geometric projection from a global model, i.e., the pretrained model, onto a sphere centered at a specialized model, i.e, a fine-tuned model. This formulation naturally extends our earlier notion of specialization, which we showed to be equivalent to computing a barycenter in the variational parameter space, toward multi-model aggregation via barycentric averaging that minimizes the average information-geometric discrepancy across multiple fine-tuned models. Our approach thus provides an interpretable and theoretically grounded merging mechanism that generalizes existing techniques such as Fisher-weighted averaging Matena and Raffel (2022), as well as mixture and product-of-experts methods.
Related Work: Model Merging.
Originally proposed in the context of FL to mitigate communication overhead and enhance privacy McMahan et al. (2017), model merging has since been adopted in both computer vision and large language models Goddard et al. (2024). Wortsman et al. (2022) demonstrated that averaging the weights of models fine-tuned under varied hyperparameters can improve both accuracy and out-of-distribution robustness. Using only a few fine-tuned models, Stock Jang et al. (2024) achieved robust merges via layer-wise linear interpolations that explicitly operate in the Euclidean geometry of the parameter space. In contrast, our method focuses on the manifold geometry of the variational posteriors.
Generalization via $D$-Barycenters.
Let $\{p_{k}\}_{k=1}^{N}$ denote the variational posteriors (e.g., IVON-trained posteriors of fine-tuned FMs), and let $p_{k}(y|x)$ denote the predictive distribution of the $k^{\text{th}}$ FM.
  • Forward KL. With $D=\mathrm{KL}(p_{k}\|q)$, the minimizer is the mixture in posterior space: $p_{D}^{\star}(\theta)=\sum_{k=1}^{N}w_{k}\,p_{k}(\theta)$. After marginalizing over $\theta$, the resulting predictive distribution also mixes pointwise as $p_{D}^{\star}(y\mid x)=\sum_{k=1}^{N}w_{k}\,p_{k}(y\mid x)$, a construction commonly referred to as the Mixture of Experts.
  • Reverse KL. With $D=\mathrm{KL}(q\|p_{k})$, the solution is the log-opinion pool or Product of Experts: $p_{D}^{\star}(\theta)\propto\prod_{k=1}^{N}p_{k}(\theta)^{w_{k}}$. In exponential families, this corresponds to natural-parameter averaging. For Gaussians posteriors, $\Lambda^{\star}=\sum_{k=1}^{N}w_{k}\,\Lambda_{k}$ and $\mu^{\star}=(\Lambda^{\star})^{-1}\sum_{k=1}^{N}w_{k}\,\Lambda_{k}\mu_{k}$, where $\Lambda_{k}$ denotes the precision matrix of the $k^{\text{th}}$ model. This formulation connects directly to Fisher merging Matena and Raffel (2022).
  • Wasserstein-2. With $D=\operatorname*{W_{2}^{2}}2(p_{k}||q)$, the minimizer is the Wasserstein-2 barycenter. For Gaussians distributions, the barycenter remains Gaussian with $\Sigma_{\operatorname*{W_{2}^{2}}2}=\left(\sum_{k=1}^{N}w_{k}\Sigma_{k}^{\frac{1}{2}}\right)^{2},\quad\mu_{\operatorname*{W_{2}^{2}}2}=\sum_{k=1}^{N}w_{k}\mu_{k}$ often yielding more robust summaries than naive parameter averaging.
Practical implications.
Barycentric merging provides a single, interpretable control (the weights $\{w_{k}\}$) to balance global and domain-specific knowledge. It recovers popular FM-merging schemes as special cases and admits closed-form solutions for common variational families (e.g., diagonal Gaussians) under widely used divergences. When complemented by the IVON training regime, the weights can be instantiated from curvature estimates, such that higher-curvature models receive greater weight Daheim et al. (2023); Shen et al. (2024). Consequently, high-uncertainty (low-curvature) models are down-weighted during aggregation.
Limitations and Possible Extensions.
Like most merging methods in distribution space (Bayesian) or parameter space (deterministic), our approach assumes architecturally aligned models, i.e., compatible layers and widths, to enable layer-wise aggregation. This constraint is particularly limiting for FMs, where specialized models are often smaller than the pretrained backbone. As a next step, we aim to relax this assumption through Gromov–Wasserstein optimal transport maps (e.g., see Delon et al. (2022); Le et al. (2022)), which enable mappings between spaces of different dimensionalities. Furthermore, we plan to conduct FM-scale experiments to assess the scalability and efficiency of the method in large-scale settings.
E.2 Incremental Learning